# Transformers Provably Solve Parity Efficiently with Chain of Thought

### v1

@@ Abstract
This work provides the first theoretical analysis of training transformers to solve complex problems by recursively generating intermediate states, analogous to fine-tuning for chain-of-thought (CoT) reasoning. We consider training a one-layer transformer to solve the fundamental $k$-parity problem, extending the work on RNNs by Wies et al. (2023). We establish three key results: (1) any finite-precision gradient-based algorithm, without intermediate supervision, requires substantial iterations to solve parity with finite samples. (2) In contrast, when intermediate parities are incorporated into the loss function, our model can learn parity in one gradient update when aided by teacher forcing, where ground-truth labels of the reasoning chain are provided at each generation step. (3) Even without teacher forcing, where the model must generate CoT chains end-to-end, parity can be learned efficiently if augmented data is employed to internally verify the soundness of intermediate steps. These results rigorously show that task decomposition and stepwise reasoning naturally arise from optimizing transformers with CoT; moreover, self-consistency checking can improve reasoning ability, aligning with empirical studies of CoT.
@@ 1 Introduction
A promising approach to overcome these limitations is chain-of-thought (CoT) reasoning, where the model is prompted or fine-tuned to solve complex tasks step-by-step by explicitly making intermediate reasoning steps to arrive at the desired answers (Wei et al., 2022; Kojima et al., 2022). Since its discovery, CoT reasoning has been shown to significantly enhance the problem-solving capabilities of LLMs while also increasing the interpretability and trustworthiness of the reasoning process, and has spawned numerous prompting techniques (Liu et al., 2023; Qiao et al., 2023) and applications for a variety of downstream tasks including common-sense reasoning, mathematical problem-solving, and symbolic or multi-modal reasoning; see e.g. Zhang et al. (2023b); Yu et al. (2023); Chu et al. (2024) for surveys on CoT. In particular, besides being used as a prompting method, directly training or fine-tuning models to generate CoT chains has also been demonstrated to significantly improve multi-step reasoning performance (Nye et al., 2021; Wei et al., 2022; Zelikman et al., 2022; Lightman et al., 2024).
Despite these empirical successes, however, the theoretical understanding of the CoT mechanism and task decomposition in transformers is still limited. Existing works focus on characterizing the expressivity of transformers equipped with CoT, providing constructions which can solve certain complexity classes (Feng et al., 2023; Chiang et al., 2023; Merrill and Sabharwal, 2023, 2024; Li et al., 2024b), studying the class of functions that can be learned in-context with CoT (Li et al., 2023; Bhattamishra et al., 2024), or analyzing the estimation error of multi-step models (Hu et al., 2024). Nevertheless, such approaches do not indicate how such capabilities might emerge when training transformers to generate reasoning chains. Li et al. (2024a) analyze the training dynamics of a one-layer transformer in an in-context learning setting and show that CoT ability may be acquired; however, they do not consider explicitly training with CoT chains, which is a more difficult problem since the objective depends on the recursive application of the transformer to itself.
@@ 1.1 Related Works
A line of work aims to understand the effectiveness of CoT from the perspective of complexity theory. Feng et al. (2023) show that autoregressive transformers of constant size can solve basic arithmetic tasks by recursively generating CoT reasoning steps, which is not possible when directly generating the solution; this separation arises because looping the generated outputs back to its inputs increases the ‘effective depth’ of the model. Works such as Chiang et al. (2023); Merrill and Sabharwal (2023) study the expressivity of fixed-precision transformer architectures in terms of classes of formal languages. Merrill and Sabharwal (2024); Li et al. (2024b) show that CoT reasoning enables recognizing wider language classes, and characterizes the increased expressivity depending on the length of the reasoning chain. Sanford et al. (2024) studies the relation between transformers and massively parallel computation protocols, showing that logarithmic depth suffices to solve multi-hop induction tasks that cannot be efficiently solved by other sequence models. Additionally, Li et al. (2023); Bhattamishra et al. (2024) study the class of functions that can be learned in context by transformers with CoT from the point of view of in-context learning.
@@
The difficulty of learning parity without task decomposition is established in Shalev-Shwartz et al. (2017); Shamir (2018). The work most relevant to our paper is due to Wies et al. (2023), who study task decomposition for parity with classical Elman RNNs. They show that by incorporating intermediate states into the loss function and utilizing teacher forcing, parity can be solved in polynomial iterations and with polynomial embedding size. Theorem 5 of our paper extends this positive result to autoregressive transformers, rigorously establishing the benefits of CoT-based training.
@@ 2.1 The Parity Problem
Let $\ell_{0-1}$ be the zero-one loss. There exists an $O(e^{-d/3})$-approximate oracle $\widetilde{\nabla}$ such that [^2] The original paper states that $\mathcal{A}$ can be any iterative gradient-based algorithm which receives an $\Omega(e^{-d/3})$-approximation of the gradient at each step. However, the result is only valid for certain adversarial perturbation schemes. the output $\mathcal{A}(\theta)$ of any iterative algorithm which sequentially makes at most $O(\operatorname{poly}(d))$ queries to ~$\widetilde{\nabla}\bar{L}$ must satisfy
$\mathbb{E}_{\bm{x}}\left[\ell_{0-1}(p(\bm{x}),f_{\mathcal{A}(\theta)}(\bm{x}))\right]\geq\textstyle\frac{1}{2}-O(e^{-d})$
@@ 2.2 Transformer Model
This type of reparametrization is common in the literature (Zhang et al., 2023a; Huang et al., 2023b; Mahankali et al., 2023; Kim and Suzuki, 2024).
@@
$\operatorname{TF}^{(v)}(\bm{x}_{1},\cdots,\bm{x}_{d},\bm{0}_{n},\cdots,\bm{0}_{n};\mathbf{W})=(\hat{\bm{x}}_{1},\cdots,\hat{\bm{x}}_{d+k-1})$
@@ 3.1 Hardness of Parity without CoT
where $\lVert\cdot\rVert_{n}$ is the empirical norm. It is important that the model $f_{\theta}$ is applied to each $\bm{x}^{i}$ separately and does not cross-reference between different samples, as there exist more efficient parity-learning algorithms if the data is allowed to be manipulated freely. For example, Gaussian elimination can solve parity with $O(d)$ samples and $O(d^{3})$ iterations (Raz, 2018). Moreover, this implies that neural networks trained with stochastic gradient descent can also solve parity in polynomial time (Abbe and Sandon, 2020). Instead, in our setting the model is forced to learn from the averaged gradient signal and can only implicitly utilize the correlation between samples.
@@
    If $n=e^{\Omega(d)}$ and $f_{\theta}$ has polynomially bounded gradients, there exists an $e^{-\Omega(d)}$-approximate gradient oracle $\widetilde{\nabla}$ such that with probability $1-e^{-\Omega(d)}$ over random sampling, the output $\mathcal{A}(\theta)$ of any iterative (possibly randomized) algorithm which makes at most $O(\operatorname{poly}(d))$ queries to ~$\widetilde{\nabla}L_{n}$ has $L_{2}$-loss lower bounded as
    $\mathbb{E}_{p\in P,\bm{x}}\left[(p(\bm{x})-f_{\mathcal{A}(\theta)}(\bm{x}))^{2}\right]\geq 1-e^{-\Omega(d)}.$
@@
    If $n=\Omega(d^{\nu})$ and $\lVert\nabla f_{\theta}\rVert=O(d^{\nu_{1}})$, there exists an $O(d^{-\nu_{2}})$-approximate gradient oracle $\widetilde{\nabla}$ such that with probability $1-e^{-\Omega(d)}$ over random sampling, the output $\mathcal{A}(\theta)$ of any iterative (possibly randomized) algorithm which makes at most $O(d^{\nu_{3}})$ queries to ~$\widetilde{\nabla}L_{n}$ has $L_{2}$-loss lower bounded, where $\nu=4\nu_{1}+4\nu_{2}+2\nu_{3}+2\nu_{4}+1$, as
    $\mathbb{E}_{p\in P,\bm{x}}\left[(p(\bm{x})-f_{\mathcal{A}(\theta)}(\bm{x}))^{2}\right]\geq 1-O(d^{-\nu_{4}}).$
@@
Since all $\binom{d}{k}=e^{\Theta(d)}$ parities in $P$ are pairwise orthogonal with respect to the uniform distribution $\operatorname{Unif}(\{\pm 1\}^{d})$, it follows that the variance of $\nabla\bar{L}$ is exponentially small and the target signal can be drowned out by a correspondingly small noise from the oracle. However, this is not true for the empirical distribution which cannot distinguish all elements in $P$ with only $\operatorname{poly}(d)$ samples; the empirical correlation of two random parities will generally be $\Theta(n^{-1/2})$. Therefore a more careful decorrelation argument is needed, resulting in the weaker guarantees of Theorem 2(2). Another technical difference is that Theorem 1 only considers the strong zero-one loss (more formally, their results can be seen to hold for any parity estimator $\hat{p}_{\mathcal{A}(\theta)}\in P$ depending on the algorithm output), while we prove the $L_{2}$ lower bound for any real-valued estimator $f_{\mathcal{A}(\theta)}$.
@@ 3.2 CoT with Teacher Forcing
When training with teacher forcing, at each position $d+1\leq m\leq d+k-1$, the ground-truth labels of the preceding intermediate states $\bm{x}_{1},\cdots,\bm{x}_{m-1}$ are fed into the transformer input to obtain the predictor $\hat{\bm{x}}_{m}$ at the $m$th position; $\hat{\bm{x}}_{m}=\operatorname{TF}(\bm{x}_{1},\cdots,\bm{x}_{m-1},\bm{0}_{n},\cdots,\bm{0}_{n};\mathbf{W})_{m}$. The loss function then computes the squared error over all states,
@@
At inference time, test inputs $\bm{x}_{1},\cdots,\bm{x}_{d}$ are randomly generated and the prediction for $\bm{y}_{\operatorname{test}}=p(\bm{x}_{1},\cdots,\bm{x}_{d})$ is computed by iterating $\operatorname{TF}$ to generate all $k$ reasoning steps without reference to ground-truth labels; $\hat{\bm{y}}_{\operatorname{test}}=\operatorname{TF}^{(k-1)}(\bm{x}_{1},\cdots,\bm{x}_{d},\bm{0}_{n},\cdots,\bm{0}_{n};\mathbf{W})_{d+k-1}$. Our positive learning result in this setting is as follows.
@@
In the above equation, if $\alpha,\beta$ are the two child nodes of $m$, the parity $x_{\alpha}x_{\beta}x_{m}\equiv 1$ will be trivial and $\left\langle\bm{x}_{m},\bm{x}_{\alpha},\bm{x}_{\beta}\right\rangle=n$. On the other hand, for nontrivial parities the interaction strength will generally be $O(\sqrt{n\log d})$ due to sample concentration. For sufficiently large $n$, the trivial parities dominate, allowing us to extract the leading term. Performing these computations up to fourth order interaction terms, we can show that the dominating signal of the gradient is $\Theta(d^{-2})$ when $j=\mathsf{c}_{1}[m],\mathsf{c}_{2}[m]$ and $O(d^{-2-\epsilon/8})$ otherwise. Hence the transformer learns to increase only the weights at the relevant positions for each subtask, and is able to compute the desired $2$-parity $\hat{\bm{x}}_{m}\approx\phi(\frac{1}{2}(\hat{\bm{x}}_{\mathsf{c}_{1}[m]}+\hat{\bm{x}}_{\mathsf{c}_{2}[m]}))\approx\hat{\bm{x}}_{\mathsf{c}_{1}[m]}\hat{\bm{x}}_{\mathsf{c}_{2}[m]}$ during its forward pass. ∎
@@ 3.3 CoT without Teacher Forcing
In this section, we extend Theorem 5 to training a transformer without teacher forcing. The main difficulty in this setting is that wrong answers propagate to later generation steps, exponentially amplifying errors and drowning out the main gradient signals. Error amplification is also a central practical issue of CoT (Zhang and Parkes, 2023; Wang et al., 2023). To solve this issue, we make some modifications to our transformer model.
@@
Second, we implement a simple data augmentation technique where randomly generated $d$-bit strings $\bm{u}^{i}\sim\operatorname{Unif}\{\pm 1\}^{d}$ for $i\in[n^{\prime}]$ are appended to the original dataset $(\bm{x}^{i})_{i\in[n]}$. The resulting extended token embeddings are denoted as $\bm{x}_{j}^{+}=(\bm{x}_{j}^{\top}\,\bm{u}_{j}^{\top})^{\top}\in\operatorname{\mathbb{R}}^{n+n^{\prime}}$ so that $\bm{p}_{j}=((\bm{x}_{j}^{+})^{\top}\,\bm{e}_{j}^{\top})^{\top}$, and the key, query and value matrices are appropriately enlarged. The ground truth labels as well as the intermediate states for the augmented data are unknown, so they are not included in the loss function; nevertheless, unlabeled data can still suffice for self-consistency (Huang et al., 2023a). Instead, their purpose is to filter for ‘faulty reasoning’ in the following sense. If the input weights are not sufficiently trained, the output of a node $x_{j}$ will consist of all nearly $-1$s and thus be uninformative for computing any parities. If the augmented tokens newly generated in the previous iteration of $\operatorname{TF}(\cdot)$ (i.e. up to $\bm{u}_{d_{\ell-1}}$) are uninformative, we zero out the generated output on the basis that all subsequent reasoning will be wrong. This is achieved by adding the following filter after the feedforward layer $\phi$:
$\iota_{\ell}(\bm{z}^{+})=\begin{cases}\bm{0}&\lVert\bm{u}_{j}+\bm{1}_{n^{\prime}}\rVert_{\infty}<\varepsilon_{0}\;\;\text{for any}\;\;d_{\ell-2}<j\leq d_{\ell-1},\\ \bm{z}^{+}&\text{otherwise.}\end{cases}$
@@
$\textstyle\hat{\bm{x}}_{m}^{+}=\iota_{\mathsf{h}[m]}\circ\phi(\hat{\bm{z}}_{m}^{+}),\quad\hat{\bm{z}}_{m}^{+}=\sum_{j=1}^{d_{\mathsf{h}[m]-1}}\sigma_{j}(\bm{w}_{m})\hat{\bm{x}}_{j}^{+}.$ (3)
The loss is computed against the ground-truth labels as in (2). Similarly to the teacher forcing setup, each sequence of generated states depends only on each sample $\bm{x}^{i}$ and the augmented data $\mathbf{U}=(\bm{u}^{i})_{i\in[n^{\prime}]}$, so we may express
@@
for a fixed mapping $f^{\times}$, so that the samples are again not cross-referenced. By considering the propagation of gradients up the reasoning chain, the gradient of $f^{\times}$ can be shown to be bounded as:
@@
The exact exponent depends on the shape of $\phi$. Since $\phi(0)=-1$ and $\phi(1)=1$, it must hold that $\lVert\phi^{\prime}\rVert_{\infty}>2$. Conversely, any such $\lVert\phi^{\prime}\rVert_{\infty}$ may be achieved by taking $\phi$ to be locally quadratic around $0,\pm 1$ and smoothly joining the curve segments with straight lines of slope ±$\pm(2+\epsilon)$. Furthermore, such a link function can be realized by a simple shallow feedforward layer using e.g. $O(1)$ ReQU neurons. Hence the exponent $g$ can be taken to be arbitrarily close to $1.5$.
Finally, we implement a simple weight quantization method by rounding each component of $\mathbf{W}$ to the nearest integer after every update; $\mathbf{W}^{(t+1)}=\mathsf{r}[\mathbf{W}^{(t)}-\eta\widetilde{\nabla}_{\mathbf{W}}L(\mathbf{W}^{(t)},\mathbf{U})]$, where $\mathsf{r}:\operatorname{\mathbb{R}}\to\operatorname{\mathbb{Z}}$ is the nearest-integer operator. Equivalently, the gradient updates themselves are quantized. Integer-based quantization methods are widely used in practice to accelerate training and reduce memory usage (Wu et al., 2020; Jacob et al., 2018), and have been successfully implemented in LLMs to facilitate much more efficient finetuning (Dettmers et al., 2022, 2023). In our theoretical setting, quantization will also allow us to simplify computations involving propagation of information.
@@
Suppose $n=\Omega(d^{2+\epsilon})$ for $\epsilon>0$, $n^{\prime}=\operatorname{poly}(d)$, $d$ is sufficiently large and let $\widetilde{\nabla}$ be any $O(d^{-2-\epsilon/8})$-approximate gradient oracle. Set $\mathbf{W}^{(0)}=\bm{0}$ and $\eta=\Theta(d^{2+\epsilon/16})$. Then for any target parity $p\in P$, it holds with probability $1-\exp(-d^{(\epsilon\wedge 1)/2})$ over random sampling of (original and augmented) data that the sequence of updates $\mathbf{W}^{(t+1)}=\mathsf{r}[\mathbf{W}^{(t)}-\eta\widetilde{\nabla}L(\mathbf{W}^{(t)})]$ w.r.t. the objective (4) without teacher forcing achieves exponentially small loss $\lVert\hat{\bm{y}}_{\operatorname{test}}-\bm{y}_{\operatorname{test}}\rVert_{\infty}\leq\exp(-\Omega(d^{\epsilon/16}))$ in $\log_{2}k$ iterations.

### v2

@@ Abstract
This work provides the first theoretical analysis of training transformers to solve complex problems by recursively generating intermediate states, analogous to fine-tuning for chain-of-thought (CoT) reasoning. We consider training a one-layer transformer to solve the fundamental $k$-parity problem, extending the work on RNNs by Wies et al. (2023). We establish three key results: (1) any finite-precision gradient-based algorithm, without intermediate supervision, requires substantial iterations to solve parity with finite samples. (2) In contrast, when intermediate parities are incorporated into the loss function, our model can learn parity in one gradient update when aided by teacher forcing, where ground-truth labels of the reasoning chain are provided at each generation step. (3) Even without teacher forcing, where the model must generate CoT chains end-to-end, parity can be learned efficiently if augmented data is employed to internally verify the soundness of intermediate steps. Our findings, supported by numerical experiments, show that task decomposition and stepwise reasoning naturally arise from optimizing transformers with CoT; moreover, self-consistency checking can improve multi-step reasoning ability, aligning with empirical studies of CoT.
@@ 1 Introduction
A promising approach to overcome these limitations is chain-of-thought (CoT) reasoning, where the model is prompted or fine-tuned to solve complex tasks step-by-step by explicitly making intermediate reasoning steps to arrive at the desired answers (Wei et al., 2022; Kojima et al., 2022). Since its discovery, CoT reasoning has been shown to significantly enhance the problem-solving capabilities of LLMs while also increasing the interpretability and trustworthiness of the reasoning process, and has spawned numerous prompting techniques (Liu et al., 2023; Qiao et al., 2023) and applications for a variety of downstream tasks including common-sense reasoning, mathematical problem-solving, and symbolic or multi-modal reasoning; see e.g. Zhang et al. (2023b); Yu et al. (2023); Chu et al. (2024) for surveys on CoT. In particular, besides being used as a prompting method, directly training or fine-tuning models to generate CoT has also been shown to significantly improve multi-step reasoning performance (Nye et al., 2021; Wei et al., 2022; Zelikman et al., 2022; Lightman et al., 2024).
Despite these empirical successes, however, the theoretical understanding of the CoT mechanism and task decomposition in transformers is still limited. Existing works focus on characterizing the expressivity of transformers equipped with CoT, providing constructions which can solve certain complexity classes (Feng et al., 2023; Chiang et al., 2023; Merrill & Sabharwal, 2023; 2024; Li et al., 2024b), studying the class of functions that can be learned in-context with CoT (Li et al., 2023; Bhattamishra et al., 2024), or analyzing the estimation error of multi-step models (Hu et al., 2024). Nevertheless, such approaches do not indicate how such capabilities might emerge when training transformers to generate reasoning chains. Li et al. (2024a) analyze the training dynamics of a one-layer transformer in an in-context learning setting and show that CoT ability may be acquired; however, they do not consider explicitly training with CoT chains, which is a more difficult problem since the objective depends on the recursive application of the transformer to itself.
@@
  • We conduct numerical experiments supporting our findings (Section 4 and Appendix D).
@@ 1.1 Related Works
A line of work aims to understand the effectiveness of CoT from the perspective of complexity theory. Feng et al. (2023) show that autoregressive transformers of constant size can solve basic arithmetic tasks by recursively generating CoT reasoning steps, which is not possible when directly generating the solution; this separation arises because looping the generated outputs back to its inputs increases the ‘effective depth’ of the model. Works such as Chiang et al. (2023); Merrill & Sabharwal (2023) study the expressivity of fixed-precision transformer architectures in terms of classes of formal languages. Merrill & Sabharwal (2024); Li et al. (2024b) show that CoT reasoning enables recognizing wider language classes, and characterizes the increased expressivity depending on the length of the reasoning chain. Sanford et al. (2024) studies the relation between transformers and massively parallel computation protocols, showing that logarithmic depth suffices to solve multi-hop induction tasks that cannot be efficiently solved by other sequence models. Additionally, Li et al. (2023); Bhattamishra et al. (2024) study the class of functions that can be learned in context by transformers with CoT from the point of view of in-context learning.
@@
The difficulty of learning parity without task decomposition is established in Shalev-Shwartz et al. (2017); Shamir (2018). The work most relevant to our paper is Wies et al. (2023), which study task decomposition for parity with classical Elman RNNs. They show that by incorporating intermediate states into the loss function and utilizing teacher forcing, parity can be solved with polynomial iterations and embedding size. Our Theorem 5 extends this positive result to autoregressive transformers, rigorously establishing the benefits of CoT-based training.
@@ 2.1 The Parity Problem
Let $\ell_{0-1}$ be the zero-one loss. There exists an $O(e^{-d/3})$-approximate oracle $\widetilde{\nabla}$ such that [^2] The original paper states that $\mathcal{A}$ can be any iterative gradient-based algorithm which receives an $\Omega(e^{-d/3})$-approximation of the gradient at each step. However, the result is only valid for certain adversarial perturbation schemes. the output $\theta(\mathcal{A})$ of any iterative algorithm $\mathcal{A}$ which sequentially makes at most $O(\operatorname{poly}(d))$ queries to ~$\widetilde{\nabla}\bar{L}$ must satisfy
$\mathbb{E}_{\bm{x}}\left[\ell_{0-1}(p(\bm{x}),f_{\theta(\mathcal{A})}(\bm{x}))\right]\geq\textstyle\frac{1}{2}-O(e^{-d})$
@@ 2.2 Transformer Model
This type of reparametrization is common in the literature to make dynamical analysis tractable (Zhang et al., 2023a; Huang et al., 2023b; Mahankali et al., 2023; Kim & Suzuki, 2024).
@@
$\operatorname{TF}^{(k-1)}(\bm{x}_{1},\cdots,\bm{x}_{d},\bm{0}_{n},\cdots,\bm{0}_{n};\mathbf{W})=(\hat{\bm{x}}_{1},\cdots,\hat{\bm{x}}_{d+k-1})$
@@ 3.1 Hardness of Parity without CoT
where $\lVert\cdot\rVert_{n}$ is the empirical norm. It is important that the model $f_{\theta}$ is applied to each $\bm{x}^{i}$ separately and does not cross-reference between different samples, as there exist more efficient parity-learning algorithms if the data is allowed to be manipulated freely. For example, Gaussian elimination can solve parity with $O(d)$ samples and $O(d^{3})$ iterations (Raz, 2018). Moreover, this implies that neural networks trained with stochastic gradient descent can also solve parity in polynomial time (Abbe & Sandon, 2020). Instead, in our setting the model is forced to learn from the averaged gradient signal and can only implicitly utilize the correlation between samples.
@@
    If $n=e^{\Omega(d)}$ and $f_{\theta}$ has polynomially bounded gradients, there exists an $e^{-\Omega(d)}$-approximate gradient oracle $\widetilde{\nabla}$ such that with probability $1-e^{-\Omega(d)}$ over random sampling, the output $\theta(\mathcal{A})$ of any iterative (possibly randomized) algorithm which makes at most $O(\operatorname{poly}(d))$ queries to ~$\widetilde{\nabla}L_{n}$ has $L_{2}$-loss lower bounded as
    $\mathbb{E}_{p\in P,\bm{x}}\left[(p(\bm{x})-f_{\theta(\mathcal{A})}(\bm{x}))^{2}\right]\geq 1-e^{-\Omega(d)}.$
@@
    If $n=\Omega(d^{\nu})$ and $\lVert\nabla f_{\theta}\rVert=O(d^{\nu_{1}})$, there exists an $O(d^{-\nu_{2}})$-approximate gradient oracle $\widetilde{\nabla}$ such that with probability $1-e^{-\Omega(d)}$ over random sampling, the output $\theta(\mathcal{A})$ of any iterative (possibly randomized) algorithm which makes at most $O(d^{\nu_{3}})$ queries to ~$\widetilde{\nabla}L_{n}$ has $L_{2}$-loss lower bounded, where $\nu=4\nu_{1}+4\nu_{2}+2\nu_{3}+2\nu_{4}+1$, as
    $\mathbb{E}_{p\in P,\bm{x}}\left[(p(\bm{x})-f_{\theta(\mathcal{A})}(\bm{x}))^{2}\right]\geq 1-O(d^{-\nu_{4}}).$
@@
Since all $\binom{d}{k}=e^{\Theta(d)}$ parities in $P$ are pairwise orthogonal with respect to the uniform distribution $\operatorname{Unif}(\{\pm 1\}^{d})$, it follows that the variance of $\nabla\bar{L}$ is exponentially small and the target signal can be drowned out by a correspondingly small noise from the oracle. However, this is not true for the empirical distribution which cannot distinguish all elements in $P$ with only $\operatorname{poly}(d)$ samples; the empirical correlation of two random parities will generally be $\Theta(n^{-1/2})$. Therefore a more careful decorrelation argument is needed, resulting in the weaker guarantees of Theorem 2(2). Another technical difference is that Theorem 1 only considers the strong zero-one loss (more formally, their results can be seen to hold for any parity estimator $\hat{p}_{\theta(\mathcal{A})}\in P$ depending on the algorithm output), while we prove the $L_{2}$ lower bound for any real-valued estimator $f_{\theta(\mathcal{A})}$.
@@ 3.2 CoT with Teacher Forcing
When training with teacher forcing, at each position $d+1\leq m\leq d+k-1$, the ground-truth labels of the preceding intermediate states $\bm{x}_{1},\cdots,\bm{x}_{m-1}$ are fed into the transformer input to obtain the predictor $\hat{\bm{x}}_{m}$ at the $m$th position,
$\hat{\bm{x}}_{m}=\operatorname{TF}(\bm{x}_{1},\cdots,\bm{x}_{m-1},\bm{0}_{n},\cdots,\bm{0}_{n};\mathbf{W})_{m}.$
The loss function then computes the squared error over all states,
@@
At inference time, test inputs $\bm{x}_{1},\cdots,\bm{x}_{d}$ are randomly generated and the prediction for $\bm{y}_{\operatorname{test}}=p(\bm{x}_{1},\cdots,\bm{x}_{d})$ is computed by iterating $\operatorname{TF}$ to generate all $k-1$ reasoning steps without reference to ground-truth labels; $\hat{\bm{y}}_{\operatorname{test}}=\operatorname{TF}^{(k-1)}(\bm{x}_{1},\cdots,\bm{x}_{d},\bm{0}_{n},\cdots,\bm{0}_{n};\mathbf{W})_{d+k-1}$. Our positive learning result in this setting is as follows.
@@
In the above equation, if $\alpha,\beta$ are the two child nodes of $m$, the parity $x_{\alpha}x_{\beta}x_{m}\equiv 1$ will be trivial and $\left\langle\bm{x}_{m},\bm{x}_{\alpha},\bm{x}_{\beta}\right\rangle=n$. On the other hand, for nontrivial parities the interaction strength will generally be $O(\sqrt{n\log d})$ due to sample concentration. For sufficiently large $n$, the trivial parities dominate, allowing us to extract the leading term. Performing these computations up to fourth order interaction terms, we show that the dominating signal of the gradient is $\Theta(d^{-2})$ when $j=\mathsf{c}_{1}[m],\mathsf{c}_{2}[m]$ and $O(d^{-2-\epsilon/8})$ otherwise. Hence the transformer learns to increase only the weights at the relevant positions for each subtask, and is able to compute the desired $2$-parity $\hat{\bm{x}}_{m}\approx\phi(\frac{1}{2}(\hat{\bm{x}}_{\mathsf{c}_{1}[m]}+\hat{\bm{x}}_{\mathsf{c}_{2}[m]}))\approx\hat{\bm{x}}_{\mathsf{c}_{1}[m]}\hat{\bm{x}}_{\mathsf{c}_{2}[m]}$ at each node during its forward pass. The full proof is provided in Appendix B. ∎
@@ 3.3 CoT without Teacher Forcing
In this section, we extend Theorem 5 to training a transformer without teacher forcing. The main difficulty in this setting is that wrong answers propagate to later generation steps, exponentially amplifying errors and drowning out the main gradient signals. Error amplification or accumulation is also a central practical issue of CoT (Zhang & Parkes, 2023; Wang et al., 2023). To solve this issue, we make some modifications to our transformer model.
@@
Second, we implement a data augmentation technique where random $d$-bit strings $\bm{u}^{i}\sim\operatorname{Unif}(\{\pm 1\}^{d})$, $i\in[n^{\prime}]$ are appended to the original dataset $(\bm{x}^{i})_{i\in[n]}$. The resulting augmented tokens are denoted as $\bm{x}_{j}^{+}=(\bm{x}_{j}^{\top}\,\bm{u}_{j}^{\top})^{\top}\in\operatorname{\mathbb{R}}^{n+n^{\prime}}$, $\bm{u}_{j}=(u_{j}^{i})_{i=1}^{n^{\prime}}$ so that $\bm{p}_{j}=((\bm{x}_{j}^{+})^{\top}\,\bm{e}_{j}^{\top})^{\top}$ (the notation is extended to $j>d$), and the key, query and value matrices are appropriately enlarged. The ground truth labels as well as the intermediate states for the augmented data are unknown, so they are not included in the loss function. Nevertheless, unlabeled data can still suffice for self-consistency (Huang et al., 2023a); their purpose is to filter for ‘faulty reasoning’ in the following sense. If the weights are not sufficiently trained, the output of a node $x_{j}$ will consist of all nearly $-1$s and thus be uninformative for computing any parities. If the augmented tokens newly generated in the previous iteration of $\operatorname{TF}(\cdot)$ (i.e. up to $\bm{u}_{d_{\ell-1}}$) are uninformative, we zero out its output on the basis that all subsequent reasoning will be wrong. This is achieved by adding the following filter after the feedforward layer $\phi$:
$\forall\bm{z}^{+}\in\operatorname{\mathbb{R}}^{n+n^{\prime}},\quad\iota_{\ell}(\bm{z}^{+})=\begin{cases}\bm{0}&\lVert\bm{u}_{j}+\bm{1}_{n^{\prime}}\rVert_{\infty}<\varepsilon_{0}\;\;\text{for any}\;\;d_{\ell-2}<j\leq d_{\ell-1},\\ \bm{z}^{+}&\text{otherwise.}\end{cases}$
@@
$\hat{\bm{x}}_{m}^{+}=\iota_{\mathsf{h}[m]}\circ\phi(\hat{\bm{z}}_{m}^{+}),\quad\hat{\bm{z}}_{m}^{+}=\sum_{j=1}^{d_{\mathsf{h}[m]-1}}\sigma_{j}(\bm{w}_{m})\hat{\bm{x}}_{j}^{+}.$ (3)
The loss is computed against the ground-truth labels as in (2). As before, each sequence of generated states depends only on each sample $\bm{x}^{i}$ and the augmented data $\mathbf{U}=(\bm{u}^{i})_{i\in[n^{\prime}]}$, so we may express
@@
for a fixed mapping $f^{\times}$, so that the samples are again not cross-referenced. By considering the propagation of gradients up the chain, the gradient of $f^{\times}$ can be shown to be bounded as follows.
@@
The exact exponent $g$ depends on the shape of $\phi$. Since $\phi(0)=-1$ and $\phi(1)=1$, it must hold that $\lVert\phi^{\prime}\rVert_{\infty}>2$. Conversely, any such $\lVert\phi^{\prime}\rVert_{\infty}$ may be achieved by taking $\phi$ to be locally quadratic around $0,\pm 1$ and smoothly joining the curve segments with straight lines of slope ±$\pm(2+\epsilon)$. Furthermore, such a link function can be realized by a simple shallow feedforward layer using e.g. $O(1)$ ReQU neurons. Hence $g$ can be taken to be arbitrarily close to $1.5$.
Finally, we implement a simple weight quantization method by rounding each entry of $\mathbf{W}$ to the nearest integer after every update; $\mathbf{W}^{(t+1)}=\mathsf{r}[\mathbf{W}^{(t)}-\eta\widetilde{\nabla}_{\mathbf{W}}L(\mathbf{W}^{(t)},\mathbf{U})]$, where $\mathsf{r}:\operatorname{\mathbb{R}}\to\operatorname{\mathbb{Z}}$ is the nearest-integer operator. Equivalently, the gradients themselves are quantized. Integer-based quantization methods are widely used in practice to accelerate training and reduce memory usage (Wu et al., 2020; Jacob et al., 2018), and have been successfully implemented in LLMs to facilitate efficient fine-tuning (Dettmers et al., 2022; 2023). In our theoretical setting, quantization also allows us to simplify computations involving propagation of error.
@@
Suppose $n=\Omega(d^{2+\epsilon})$ for $\epsilon>0$, $n^{\prime}=\operatorname{poly}(d)$, [^4] Any polynomial order suffices for the number of augmented data samples. $d$ is sufficiently large and let $\widetilde{\nabla}$ be any $O(d^{-2-\epsilon/8})$-approximate gradient oracle. Set $\mathbf{W}^{(0)}=\bm{0}$ and $\eta=\Theta(d^{2+\epsilon/16})$. Then for any target parity $p\in P$, it holds with probability $1-\exp(-d^{(\epsilon\wedge 1)/2})$ over random sampling of (original and augmented) data that the sequence of updates $\mathbf{W}^{(t+1)}=\mathsf{r}[\mathbf{W}^{(t)}-\eta\widetilde{\nabla}L(\mathbf{W}^{(t)},\mathbf{U})]$ w.r.t. the objective (4) without teacher forcing achieves exponentially small loss $\lVert\hat{\bm{y}}_{\operatorname{test}}-\bm{y}_{\operatorname{test}}\rVert_{\infty}\leq\exp(-\Omega(d^{\epsilon/16}))$ in $\log_{2}k$ iterations.